# Statistical Rounding Error Analysis for the Computation of Random Vectors and Matrices

### v2

@@ Appendix D Proof of Theorem 5
$=\frac{a_{ik}\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{\hat{l}_{ij}\hat{u}_{jk}\prod_{r}^{\left(k-j+1\right)}{\left(1+\delta_{r}\right)}}}{\hat{u}_{kk}}\left(1+\delta_{o}\right)$ (71)
=$\overset{\left(b\right)}{=}\frac{a_{ik}\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{l_{ij}u_{jk}\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}}}{u_{kk}\left(1+\eta_{k}\right)}$ (72)
$\overset{\left(c\right)}{\approx}\frac{a_{ik}\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{l_{ij}u_{jk}\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}}}{u_{kk}}\left(1-\eta_{k}\right),$ (73)
where we use (63), and define $\eta_{k}=\frac{\hat{u}_{kk}-u_{kk}}{u_{kk}}$ with the mean zero and variance $\sigma_{\eta_{k}}^{2}=\frac{\mathbb{V}\left(\Delta u_{kk}\right)}{\mathbb{V}\left(u_{kk}\right)+\left[\mathbb{E}\left(u_{kk}\right)\right]^{2}}$ at $\left(b\right)$, and $\left(c\right)$ follows the Taylor expansion.
$\Delta l_{ik}$ $\approx o_{i}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right]$
$-\sum_{j=1}^{k-1}{p_{j}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]}.$
Therefore, based on Lemma 2, Lemma 5, Lemma 7 and Lemma 8, the variance of $\Delta l_{ik}$ can be derived by
$\mathbb{V}\left(\Delta l_{ik}\right)$ $\approx\mathbb{V}\left(o_{i}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right]\right)$
$+\mathbb{V}\left(\sum_{j=1}^{k-1}{p_{j}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$
$-2\sum_{j=1}^{k-1}{\mathbb{C}\left(o_{i}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right],p_{j}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
$\approx\frac{\left(m-4\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}$
$+\frac{\sum_{j=1}^{k-1}{\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}-1\right]}}{\left(m-k-1\right)\left(m-k-3\right)}$
$-2\sum_{j=1}^{k-1}{\mathbb{E}\left(o_{i}p_{j}\left[\left(1+\eta_{k}^{2}\right)\prod_{i}^{\left(k-j+1\right)}{\left(1+\delta_{i}^{2}\right)-1}\right]\right)}$
$=\frac{\left(m-4\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}$
$+\frac{\sum_{j=1}^{k-1}{\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}-1\right]}}{\left(m-k-1\right)\left(m-k-3\right)}$
@@
$=\frac{\left(m-6\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}$
$+\frac{\left(1+\sigma_{\eta_{k}}^{2}\right)\sum_{i=j}^{k-1}{\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}}-k+1}{\left(m-k-1\right)\left(m-k-3\right)}$
$-\frac{2}{\left(m-k-1\right)\left(m-k-3\right)}\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\frac{\left(1+\sigma^{2}\right)^{2}\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}-k+2\right].$ (74)
Therefore, the theorem holds.

### v3

@@ Appendix D Proof of Theorem 5
$=\frac{a_{ik}\prod_{r}^{\left(k-1\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{\hat{l}_{ij}\hat{u}_{jk}\prod_{r}^{\left(k-j+1\right)}{\left(1+\delta_{r}\right)}}}{\hat{u}_{kk}}\left(1+\delta_{o}\right)$
=$\overset{\left(b\right)}{=}\frac{a_{ik}\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{l_{ij}u_{jk}\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}}}{u_{kk}\left(1+\eta_{k}\right)}$
$\overset{\left(c\right)}{\approx}\frac{a_{ik}\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)}-\sum_{j=1}^{k-1}{l_{ij}u_{jk}\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}}}{u_{kk}}\left(1-\eta_{k}\right).$
$\mathbb{V}\left(\Delta l_{ik}\right)$ $\approx\mathbb{V}\left(o_{ik}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right]\right)+\mathbb{V}\left(\sum_{j=1}^{k-1}{p_{ijk}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]}\right)$ (71)
$-2\sum_{j=1}^{k-1}{\mathbb{C}\left(o_{ik}\left[\left(1-\eta_{k}\right)\prod_{r}^{\left(k\right)}{\left(1+\delta_{r}\right)-1}\right],p_{ijk}\left[\left(1-\eta_{k}\right)\left(1+\epsilon_{j}\right)\left(1+\eta_{j}\right)\prod_{r}^{\left(k-j+2\right)}{\left(1+\delta_{r}\right)}-1\right]\right)}$
$\approx\frac{\left(m-4\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}+\frac{\sum_{j=1}^{k-1}{\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}-1\right]}}{\left(m-k-1\right)\left(m-k-3\right)}$
$-2\sum_{j=1}^{k-1}{\mathbb{E}\left(o_{ik}p_{ijk}\left[\left(1+\eta_{k}^{2}\right)\prod_{i}^{\left(k-j+1\right)}{\left(1+\delta_{i}^{2}\right)-1}\right]\right)}$
$=\frac{\left(m-4\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}+\frac{\sum_{j=1}^{k-1}{\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}-1\right]}}{\left(m-k-1\right)\left(m-k-3\right)}$
@@
$=\frac{\left(m-6\right)\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\left(1+\sigma^{2}\right)^{k}-1\right]}{\left(m-k-1\right)\left(m-k-3\right)}+\frac{\left(1+\sigma_{\eta_{k}}^{2}\right)\sum_{i=j}^{k-1}{\left(1+\sigma_{\epsilon_{j}}^{2}\right)\left(1+\sigma_{\eta_{j}}^{2}\right)\left(1+\sigma^{2}\right)^{k-j+2}}-k+1}{\left(m-k-1\right)\left(m-k-3\right)}$
$-\frac{2}{\left(m-k-1\right)\left(m-k-3\right)}\left[\left(1+\sigma_{\eta_{k}}^{2}\right)\frac{\left(1+\sigma^{2}\right)^{2}\left[\left(1+\sigma^{2}\right)^{k-2}-1\right]}{\sigma^{2}}-k+2\right].$
Therefore, based on Lemma 1, Lemma 4, Lemma 6 and Lemma 7, the variance of $\Delta l_{ik}$ can be derived in (71) at the top of the next page. Therefore, the theorem holds.